OpenStudy (anonymous):

Determine if (-5/2, 1) is a solution to -2x+3y=2

OpenStudy (australopithecus):

sub the point into the equation does it equal to 2 on both sides?

OpenStudy (australopithecus):

-2(-5/2)+3(1)=2

OpenStudy (australopithecus):

remember any whole number can be expressed as a fraction just use 1 as the denominator (-2/1)(-5/2) +3(1) = 2 do you know how to multiply fractions?

OpenStudy (australopithecus):

its just numerator*numerator, denominator*denominator

OpenStudy (australopithecus):

so (-2*-5)/(1*2) + 3(1) = 2

OpenStudy (australopithecus):

you can do the rest if you get 2=2 then the answer is yes if not the answer is no obviously

OpenStudy (anonymous):

-2*-5=10 1*2=2 10/2=5 5+3=8 8-1=7

OpenStudy (australopithecus):

I dont know where the 10 came from

OpenStudy (australopithecus):

and your solution seems incorrect

OpenStudy (australopithecus):

please read what I wrote

OpenStudy (australopithecus):

remember that a point is (x,y)

OpenStudy (australopithecus):

well defined as

OpenStudy (australopithecus):

oh I see you split it up I dont recommend doing that you may misplace numbers

OpenStudy (australopithecus):

(-2*-5)/(1*2) + 3(1) = 2 remember a - * - = + - * + = - + * + = +

OpenStudy (australopithecus):

so we are left with (10/2) + 3 = 2

OpenStudy (anonymous):

-2*-5=-10 1*2=2 -10/2=-5 -5+3=2

OpenStudy (australopithecus):

-2*-5 =/= -10

OpenStudy (australopithecus):

- * - = + - * + = - + * + = +

OpenStudy (anonymous):

-2 sorry

OpenStudy (australopithecus):

no please review what I posted just above your last response

OpenStudy (australopithecus):

-1 * -1 = 1 -2 * -1 = 2 -3 * -2 = 6 etc -1*+1 = -1 -2 * +1 =-2 -3 * +2 = -6 etc +2 * +1 = 2 +3 * +2 = 6 etc

OpenStudy (australopithecus):

so based on this we assume by rule (-)(+) = - (-)(-) = + (+)(+) = +

OpenStudy (australopithecus):

Memorize this it is fundamental and vital to your success in mathematics

OpenStudy (anonymous):

-2*-5=10 1*2=2 10/2=5 5+3=8

OpenStudy (anonymous):

thats what i did

OpenStudy (australopithecus):

dont rearrange the equation

OpenStudy (australopithecus):

I will post the proper solution for you please follow my example in the future

OpenStudy (australopithecus):

-2(-5/2)+3(1)=2 (-2*-5)/(1*2) + 3(1) = 2 (10/2)+3(1) = 2 5 + 3 = 2 8 = 2 \[8 \neq 2\] 8 does not equal 2 Thus the point does not exist, what does this mean it means that in the graph -2x+3y=2 there is never a point (-5/2, 1) http://www.wolframalpha.com/input/?i=graph+-2x%2B3y%3D2+

OpenStudy (australopithecus):

which means the line of the graph never crosses the point (-5/2, 1)

OpenStudy (australopithecus):

please make sure you keep the equation intact in the future, its ok to do rough work to solve for a number but it is vital that you always go back to the original equation

OpenStudy (australopithecus):

but yeah hope this helps

OpenStudy (anonymous):

yes it don't thank you